Opened 10 years ago

Closed 10 years ago

Last modified 3 years ago

#254 closed flight processing (fixed)

GB07/07, flight day 191/2009, Inverclyde

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB07_07-2009_191_Inverclyde

Data arrived from ARSF via SATA disk on 17/7/09

Scientific objective: Effect on spectral reflectance by heather type and biophysical properties

Priority: unknown

PI: Tim Malthus

Sensors:

  • Eagle (29/09/2009)
  • Hawk (29/09/2009)
  • LIDAR (13/11/09)
  • Photography (07/01/2010)

Change History (28)

comment:1 Changed 10 years ago by knpa

  • Owner mggr deleted

comment:2 Changed 10 years ago by knpa

e-mailed gary to request vectors

comment:3 Changed 10 years ago by mggr

  • Component changed from ARSF to Processing: general

comment:4 Changed 10 years ago by mggr

  • Milestone set to 2009 Data processing completion

comment:5 Changed 10 years ago by emca

  • Owner set to emca

comment:6 Changed 10 years ago by mggr

Note the PI has moved out of the country. Suggest using Alasdair MacArthur @ Uni of Edinburgh instead.

comment:7 Changed 10 years ago by mggr

  • Description modified (diff)

Has lidar data too!

comment:8 Changed 10 years ago by mark1

Vector data has arrived and shape files have been created.

comment:9 Changed 10 years ago by emca

  • Owner changed from emca to chrfi

Processing complete. Delivery directory created at
~airborne/workspace/GB07_07-2009_191_Inverclyde/delivery

Passing to Chris for Delivery checking.

comment:10 Changed 10 years ago by chrfi

  • Owner changed from chrfi to emca

delivery checking complete.

passing to emca for delivery

comment:11 Changed 10 years ago by emca

Delivery sent to Alasdair MacArthur 29/09/2009.

Ready for archiving.

comment:12 Changed 10 years ago by emca

LIDAR processing complete

Delivery directory created at
~airborne/workspace/GB07_07-2009_191_Inverclyde/delivery/20091103

Awaiting delivery check

comment:13 Changed 10 years ago by mark1

  • Owner changed from emca to mark1

comment:14 Changed 10 years ago by mark1

Delivery checking LIDAR data set.

Missing one LIDAR flight line - emca + gary looking into this.

Minor edits made to readme file. Dem header and contents look sensible. ASCII files contain sensible information and loaded up OK into a GIS.

comment:15 Changed 10 years ago by mark1

  • Owner changed from mark1 to emca

will wait for reply from gary before rsync-ing data over to arsf_data

comment:16 Changed 10 years ago by emca

Raw RCD files have been processed to generate tiff files, and the tiffs files have been tagged.

comment:17 Changed 10 years ago by emca

LIDAR data sent to A. MacArthur 13/11/09

comment:18 Changed 10 years ago by emca

LIDAR and Hyperspectral data has been rsynced to
/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB07_07-2009_191_Inverclyde

comment:19 Changed 10 years ago by mark1

  • Description modified (diff)

Digital Photography has been dispatched to PI on 7th January 2010.

Still needs to be rsynced.

comment:20 Changed 10 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to knpa
  • Status changed from new to assigned

Archived by DTG (12/01/10).
Preparing for NEODC.

comment:21 Changed 10 years ago by knpa

Tarball created and NEODC informed ready for transfer.

comment:22 Changed 10 years ago by knpa

Tarball was awaiting transfer to NEODC when it turned out that the photography data had not been rsynced before the archive process was started.
Mark1 has now rsynced and is ready to be archived again.

comment:23 Changed 10 years ago by knpa

  • Description modified (diff)

comment:24 Changed 10 years ago by knpa

No webcam for this project.

comment:25 Changed 10 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Archived and backed up by NEODC.
Closing ticket.

comment:26 Changed 5 years ago by dap

Delivery checking reprocessed data for delivery to PI:

  • Copied stylesheets from line_information to doc

Delivering re-processed data to PI via FTP

comment:27 Changed 3 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:28 Changed 3 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.